Andre's Dog House had current assets of $67,200 and current liabilities of $71,100 last year. This year, the current assets are $82,600 and the current liabilities are $85,100. The depreciation expense for the past year is $9,600 and the interest paid is $8,700. What is the amount of the change in net working capital?
Answer #1

Change in Net working capital = Net working capital current year - net working capital last year

Net working capital = Current assets - Current liabilities

Net working capital current year = $82,600 - 85,100 = -$2,500

Net working capital last year = $67,200 - 71,100 = - $3,900

Change in Net working capital = -$2,500 - (- $3,900) = $1,400

Hence, amount of change in net working capital is $1,400